OpenStudy (anonymous):

What is the product of 7√10 * 3√5 simplified?

OpenStudy (anonymous):

is it 21√5 or 21√2?

OpenStudy (anonymous):

\[7 \sqrt{10}\times3\sqrt{5}\]\[(7\times3)\times \sqrt{10\times5}\]\[21\sqrt{50} = 21\times5\sqrt{2} = 105\sqrt{2}\]

OpenStudy (anonymous):

ok that is so much clearer:) thanks:)